zelma is 18 years older then her son. she is 3 times as old as him. how old are they ?



Answer by robertb(5830)

zelma = x + 18
son = x
==> x+18 = 3x
==> 18 = 2x
==> x = 9, the age of the son.
x + 18 = 27, the age of zelma.